OpenStudy (anonymous):

I think my math book is wrong...but I don't know. Could someone check and if it's not wrong, explain it to me?

OpenStudy (anonymous):

here it is

OpenStudy (nubeer):

hmm what is your math book answer?

OpenStudy (anonymous):

That is the math book answer

OpenStudy (nubeer):

hmm and what is your answer?

OpenStudy (anonymous):

I know this can't be right but I got x=2.5. I don't understand how they randomly got the extra 2 when it should be a 9.

OpenStudy (nubeer):

ok.. i also have got a different answer but not sure.. so can you put steps how you solve it?

OpenStudy (nubeer):

and it's matrix right?

OpenStudy (anonymous):

it looks like its just a typing error, try asking or calling your teacher

OpenStudy (raden):

81x - 21x = 24 60x = 24 x = 24/60 = 2/5

OpenStudy (anonymous):

Yeah it's about evaluating determinants of a matrix. How I solved the problem was: (0)(1)(3)+(7)(2)(0)+(9)(x)(9) (0)(1)(9)-(9)(2)(0)-(3)(x)(7) 0+0+81x-0-0-21x=24 60x=24 x=2.5

OpenStudy (raden):

2/5 = 4/10 = 0.4 not 2.5

OpenStudy (phi):

See http://www.wolframalpha.com/input/?i=det+%7B+%7B0%2C7%2C9%7D%2C%7B0.4%2C1%2C2%7D%2C%7B0%2C9%2C3%7D%7D x= 0.4 so there must be a typo somewheres?

OpenStudy (phi):

60x=24 x=2.5 ?? don't you think it is x= 24/60 = 2/5

OpenStudy (phi):

See http://www.wolframalpha.com/input/?i=det+%7B+%7B0%2C7%2C9%7D%2C%7B0.4%2C1%2C2%7D%2C%7B0%2C9%2C3%7D%7D x= 0.4 so there must be a typo somewheres?

OpenStudy (phi):

The book's solution would be correct if the first line of the matrix were 0 7 2 (rather than 0 7 9)